What is the energy which is absorbed by a mass weighing 65kg which receives an absorbed dose of 3 μGy?


1
Expert's answer
2021-02-25T15:03:59-0500

The energy imparted to matter from any type of radiation can be found as follows:


"D=\dfrac{E}{m},"

here, "D" is the adsorbed dose, "E" energy absorbed by material of mass "m".

Then, from this formula we can find the energy absorbed by a mass of 65 kg:


"E=Dm=3\cdot10^{-6}\ Gy\cdot\dfrac{1\ J}{1\ kg}\cdot65\ kg=1.95\cdot10^{-4}\ J."